OpenStudy (maddy1251):

Someone want to check my work for me real quick? For the functions below, find the amplitude, period, phase shift, and vertical translation of the graph for the sinusoid. y=-6cos(x+pi/4)-5 I used the equation to help; F(t)=Af(B(t)-C)+D A=Amp ; B= 2pi/B (Period) C= Phase Shift ; D=Vertical Shift. Amplitude is : -6 Period is : 2pi/x (Does this mean there is range/domain restrictions?) Phase shift: pi/4 Vertical translation: -5, it's negative so it will shift left <---- Type of reflection: Vertical? Or is it meaning negative? Minimum Value of graph: -11 Max. Value of graph: -1.

OpenStudy (maddy1251):

@Owlcoffee @ganeshie8

OpenStudy (maddy1251):

Or, will it be a horizontal reflection since of the -6?

OpenStudy (maddy1251):

And could the period possibly be 0.5?

OpenStudy (owlcoffee):

I'm in a office computer right now, but I can mention a little about this. The period (the interval in which the whole function can fit and keeps on repeating after) is calculated by, if we consider the function with form: \(y=Acos(mx+c)+B\) will be calculated by \(P=\frac{ 2 \pi }{ m }\), which in this case it is just one, so therefore the period will be \(2\pi\).

OpenStudy (maddy1251):

Okay! Thank you. That makes sense. And I am thinking since -6cos, that will result in a horizontal reflection. What about the phase shift @Owlcoffee

OpenStudy (owlcoffee):

The phase shift is the same as pelleting all the points to the left pi/4 which would mean that for any angle you'll be sustracting pi/4 to it.

OpenStudy (maddy1251):

Okay, so I had that. I was just wrong on my Period, I found out what the reflection was. I think that's all I needed, thank you @Owlcoffee !

OpenStudy (owlcoffee):

No problem!

OpenStudy (mathmale):

"Period is : 2pi/x (Does this mean there is range/domain restrictions?)" NO.

OpenStudy (maddy1251):

@mathmale Oh okay! Thank you. I was just curious. Thank you so much.